in the 2004-2005 nba regular season, the phoenix suns won two more than three times as many games as they lost. the suns played 82 games. how many wins and losses did the team have?
Answer by fractalier(6550)

Let
W = number of wins
and
L = number of losses
Now we translate the English facts into algebra...
W = 2 + 3L (two more than three times what they lost)
W + L = 82 (since there are 82 games total)
Let's substitute the first equation into the second and get
(2 + 3L) + L = 82
2 + 4L = 82
Now subtract two and get
4L = 80
Divide by four...
L = 20
Subtract from 82 games...
W = 62
They won 62 and lost 20.